OpenStudy (anonymous):

Rewrite using a negative exponent. 5/x^2

OpenStudy (anonymous):

\[5/x^2\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \frac{1}{a^m} = a^{-m}\)

ganeshie8 (ganeshie8):

\(\large \frac{5}{x^2} = 5x^{-2}\)